a.The pH of an aqueous solution of 0.218 M ammonium perchlorate, NH4ClO4 (aq), is .

This solution is acidic or basic or neutral

b.he pH of an aqueous solution of 0.500 M phenol (a weak acid), C6H5OH, is .-----
Answer #1

a)

concentration of NH4ClO4 = 0.218 M

pKb of NH3 = 4.74

this is the salt of strong acid and weak base . so pH < 7

this solution is acidic.

pH = 7 - 1/2 (pKb + log C)

      = 7 - 1/2 (4.74 + log 0.218)

pH = 4.96

b)

concentration of C6H5OH = 0.500 M

pKa = 10.0

pH = 1/2 (pKa - log C)

    = 1/2 (10.0 - log 0.500)

pH = 5.15